Tek 561A And 3A1 Fixed
On Thursday, March 29, 2018 at 10:08:12 PM UTC-4, wrote:
Just thought I'd let you know, it is finally done. I had help too. Anyway :
Wired in substitute transformer for filament of CRT.
Replace Q 260 and modify to accept silicon.
Replace Q 233 & 243, silicon worked fine.
Bypassed variable vertical gain controls.
Cleaned output level switch and tub socket on calibrator.
Basic calibration of sweep and vertical gain.
Sounds easy eh ? It took some doing to figure it out and the modification for Q 260 was all wrong at first because I am not that familiar with working in a negative supply. But someone pointed out my mistake and one piece of quick math it was good to go.
Anyway, between "A Little Help From My Friends" here, on SED and the Tek Scopes group it is working pretty much perfectly. It might need a bit more calibration but I need to obtain a couple of things first. It is however, plenty usable right now.
Yeah, working with a negative supply is not intuitive to me. I usually walk away, have a cup of coffee, come back and then it's (sorta) clear.